# BEV-SUSHI: Multi-Target Multi-Camera 3D Detection and Tracking in Bird’s-Eye View

### v2

@@ 4.1 Datasets and Implementations
The original BEVFormer is developed for autonomous driving applications with moving cameras but fixed camera settings, i.e., a fixed number of cameras and fixed relative camera placement. For MTMC environments, we usually have dynamic camera systems across different scenes. AICity’24 dataset, for example, has 6 different indoor environments with cameras ranging from 7 to 16. Therefore, we re-design the spatial cross-attention layers to enable a dynamic number of cameras for different scenes and remove camera embeddings. Besides, we apply a transformation on the original 3D world coordinates to shift the origin to the center of the scene (see Sec. A.2). Moreover, we implement an additional Circle Non-Maximum Suppression (NMS) with a threshold of 0.2m to filter nearby false positives.
@@ 4.2 Results on AICity’24 Dataset
Compared with other methods on the AICity’24 benchmark leaderboard in Tab. 2, our proposed BEV-SUSHI outperforms significantly in all aspects, with the following improvements from the second place, $+9.28$ on HOTA, $+2.91$ on DetA, $+4.38$ on AssA, and $+1.85$ on LocA. The qualitative results on the AICity’24 dataset are shown in Fig. 6. We project 3D detections with track IDs to each camera view and plot their 3D locations on the floor plans. It shows that BEV-SUSHI achieves robust MTMC detection and tracking accuracies in various indoor environments.
@@ 4.3 Results on WildTrack Dataset
BEV-SUSHI (Ours) 93.4 87.5 94.3 90.2 2.4
BEV-SUSHI (Ours) 95.6 92.6 93.7 80.5 7.3
@@
As illustrated in Tab. 3, our BEV-SUSHI framework achieves state-of-the-art for IDF1 with a $+1.1$ improvement over EarlyBird [29]. However, we get a slightly lower MOTA compared to EarlyBird. This is due to the limited training data (360 frames) provided by the WildTrack dataset. Transformer-based networks, like BEVFormer, generally require larger amounts of data to unfold their potential. To demonstrate the effectiveness of BEV-SUSHI’s ReID and tracking components, we replace our detections with the ones from EarlyBird (based on MVDet [11] with ResNet-18 backbone). This version ($\dagger$) yields significant improvements on both IDF1 ($+3.3$) and MOTA ($+3.1$). Qualitative results on WildTrack are shown in Fig. 7.
Besides, we pre-train BEVFormer on the large-scale synthetic dataset (AICity’24) and finetune on the small WildTrack dataset. The results are analyzed in Sec. A.3.
@@ 4.4 Ablation Studies
To verify we have the best BEVFormer model for multi-view 3D object detection, we conduct several ablation studies on backbone selection and parameter tuning to improve the detection results. The results are shown in Tab. 4. We use a customized validation set from the AICity’24 dataset, including around 3k frames from warehouse, hospital, and retail store scenes. More results shown in Sec. A.1.
@@
The length of sequences and occlusion gaps common in MTMC benchmarks without full camera coverage of the environment pose unique challenges to our tracking model and inference. To process arbitrary long sequences, BEV-SUSHI infers windows of frames with a GNN model and connects these in time in a near-online sliding window fashion. In Tab. 5, we demonstrate how increasing the window size gradually improves the association performance, i.e., AssA. Larger windows shift responsibility from the heuristic matching between windows to the trained model, which allows us to bridge longer occlusion gaps. The last row replaces all remaining heuristics with our novel global merging block boosting overall performance by $+4.42$ HOTA.
To further illustrate the superiority of our learned GNN approach, we evaluate progressively longer subsets of a selected AICity’24 scene in Tab. 6. The BEV-KF baseline represents heuristic online trackers like EarlyBird [29] commonly only evaluated on short tracking challenges like WildTrack. From 1,000 frames to the full AICity’24 length, BEV-SUSHI only drops by $4.58$ HOTA points compared to the massive performance decrease of $43.35$ utilizing only a Kalman filter. Such trackers fail to tackle the long-term tracking challenges common in MTMC benchmarks.
Win Size Global HOTA DetA AssA LocA
480 51.87 86.12 31.26 95.55
960 62.50 86.36 45.25 95.57
1920 71.88 86.47 59.78 95.58
3840 76.80 86.49 68.22 95.58
3840 81.22 86.94 76.19 95.67
Table 5: SUSHI inference ablation on the AICity’24 test set. Rows 1 to 4 rely on heuristic matching to track overlapping sliding windows inferred by the SUSHI GNN hierarchy. Our final tracking solution (the last row) applies a global merging block to associate past tracks with incoming frames without graph overlaps. The window stride is half of the window size for all experiments.
Method # of frames HOTA DetA AssA LocA
BEV-SUSHI 1,000 95.09 94.41 95.77 97.51
3,000 93.62 94.89 92.37 97.53
5,000 91.85 94.98 88.82 97.50
10,000 91.06 93.89 88.33 97.22
23,994 90.51 93.03 88.05 97.16
BEV-KF (baseline) 1,000 63.49 89.01 45.29 97.64
3,000 45.28 89.71 22.86 97.46
5,000 35.54 88.79 14.23 97.30
10,000 27.91 87.80 8.87 97.19
23,994 20.15 86.91 4.67 97.11
@@ 5 Conclusion
This paper proposed BEV-SUSHI, an accurate and robust multi-camera 3D detection and tracking framework, based on early multi-view aggregation, in environments monitored by static multi-camera systems. The proposed framework has impressive generalizability for diverse scenes and camera settings. It also has a powerful capability for long-term association to track objects in very long videos. As a result, BEV-SUSHI achieves SOTA on both the AICity’24 dataset and the WildTrack dataset.
@@ B.3 Improvements with Detection Association
We compared the tracking performance of BEV-SUSHI with and without the proposed 2D-3D detection association algorithm in Tab. 9. The baseline result is based on the ReID features extracted from the large projected 3D bounding boxes shown in Fig. 4. With the noisy background or other objects included in the image crops, ReID feature quality will be significantly affected.
@@ Appendix C ReID Feature Quality Analysis
We evaluated the ReID feature quality by the mean average precision (mAP), rank-1, rank-5, and rank-10 accuracies. The evaluation results are shown in Tab. 10. We found that the feature quality on the WildTrack dataset is worse than that on the AICity’24 dataset. This is because i) WildTrack is a real-world dataset with more noises and diverse illuminations from different camera views; ii) 2D bounding box annotations are not as accurate as the synthetic AICity’24 dataset. Nevertheless, our BEV-SUSHI achieved impressive results on WildTrack based on these ReID features.

### v3

@@ 4.2 Implementation Details
The original BEVFormer is developed for autonomous driving applications with moving cameras but fixed camera settings, i.e., a fixed number of cameras and fixed relative camera placement. For MTMC environments, we usually have dynamic camera systems across different scenes. AICity’24 dataset, for example, has 6 different indoor environments with cameras ranging from 7 to 16. Therefore, we re-design the spatial cross-attention layers to enable a dynamic number of cameras for different scenes and remove camera embeddings. Besides, we apply a transformation on the original 3D world coordinates to shift the origin to the center of the scene (see Sec. A.2). Moreover, we implement an additional Circle Non-Maximum Suppression (NMS) with a threshold of 0.2m to filter nearby false positives. Since there is no 3D bounding box annotation in the WildTrack dataset, we remove the bounding box dimension and orientation regression losses during BEVFormer training. We use a default 3D bounding box dimension (width, length, height) of $[0.6,0.6,1.7]$ and zero rotation for later 2D-3D detection association.
2D detector.
We use DINO [41] with FAN-small backbone [45]. The detector is pre-trained on a subset (800k images) of the OpenImages dataset [17], pseudo-labeled by a DINO detector trained on 80 COCO classes. Then, the detector is fine-tuned on a proprietary dataset with more than 1.5 million images and more than 27 million objects for the person class.
ReID feature extraction.
We extract ReID features for each 2D bounding box from the 2D detector. Our person-centric ReID model is based on the SOLIDER [7], a self-supervised learning framework with a Swin-Tiny Transformer [19] backbone. The model uses image crops of size 256$\times$128 as input and outputs a feature of dimension 256 enhancing memory efficiency and throughput. We pre-train the model on a proprietary dataset of 3 million unlabeled image crops. The supervised fine-tuning is performed on a collection of both real and synthetic datasets which includes characters from Market-1501 [44], AICity’24 [32] and additional propitiatory datasets totaling the object count to 3,826 identities and image count to 82k object crops. We analyze the ReID feature quality in Appendix C.
@@ 4.3 Results on AICity’24 Dataset
MCBLT reaches SOTA on the AICity’24 benchmark on all metrics with improvements of $+9.28$ on HOTA, $+2.91$ on DetA, $+4.38$ on AssA, and $+1.85$ on LocA. These significant improvements demonstrate MCBLT’s effectiveness in advancing multi-camera tracking performance. We present qualitative results on the AICity’24 dataset in Fig. 5 by projecting 3D detections with track IDs to each camera view and plot their 3D locations on the floor plans. Overall, our quantitative and qualitative results show that MCBLT achieves robust MTMC detection and tracking accuracies in various indoor environments.
Besides, we evaluate the 2D-3D detection association algorithm to ensure we have high-quality ReID features assigned to each 3D detection. This evaluation is based on the association results between ground truth 2D and 3D annotations in the AICity’24 training set. We introduce the evaluation metric, detection association accuracy, defined as the number of corrected matches divided by the number of matched ground truth 2D bounding boxes. The evaluation results are shown in Tab. 2. We can find that even in crowded hospital scenes, our algorithm can achieve above 90% accuracy, illustrating its robustness. More visualizations based on predictions are shown in Sec. B.2.
Scene Warehouse Retail store Hospital Overall
Accuracy 99.4% 95.2% 91.2% 96.9%
@@ 4.4 Results on WildTrack Dataset
MCBLT (Ours) 93.4 87.5 94.3 90.2 2.4
MCBLT (Ours) 95.6 92.6 93.7 80.5 7.3
@@
WildTrack only provides limited training data (360 frames), while transformer-based networks, e.g., BEVFormer, generally require larger amounts of data to unfold their potential. Thus, to have a fair comparison, we report our results under two settings: (i) MCBLT with detections from BEVFormer, (ii) MCBLT using the same detections as EarlyBird (based on MVDet [11] with ResNet-18 backbone). MCBLT achieves the state-of-the-art with $+1.1$ IDF1 improvement over EarlyBird [30] even if it reaches a slightly lower MOTA score (as expected due to the transformer backbone). MCBLT significantly outperforms EarlyBird with $+3.3$ IDF1 and $+3.1$ MOTA while using the same detections, demonstrating the efficacy of our methods.
We provide qualitative results on WildTrack in Fig. 6. In addition, we present additional results on WildTrack by pre-training BEVFormer on the large-scale synthetic dataset (AICity’24) and finetuning on the small WildTrack dataset in the supplementary material Sec. A.3.
@@ 4.5 Ablation Studies
To verify we have the best BEVFormer model for multi-view 3D object detection, we conduct several ablation studies on backbone selection and parameter tuning to improve the detection results. The results are shown in Tab. 4. We use a customized validation set from the AICity’24 dataset, including around 3k frames from warehouse, hospital, and retail store scenes. More results shown in Sec. A.1. According to the results in Tab. 4, we use ResNet-101 with higher BEV resolution as our 3D detector.
@@
The length of sequences and occlusion gaps common in MTMC benchmarks without full camera coverage of the environment pose unique challenges to our tracking model and inference. To process arbitrarily long sequences, SUSHI [5] uses heuristics to stitch overlapping graphs in a sliding window fashion. In Tab. 5, we compare the heuristic matching of SUSHI with our global block for long-term association. Rows 1-4 demonstrate how increasing the window size gradually improves the association performance, i.e., HOTA and AssA. The last row replaces all heuristics with our novel global merging block while using the same learnable weights as row 4. Our global merging block increases overall performance by $+4.42$ HOTA by scaling the association window to the full length of the video.
To further illustrate the superiority of our learned GNN approach, we evaluate progressively longer subsets of a selected AICity’24 scene in Tab. 6. The BEV-KF baseline represents heuristic online trackers like EarlyBird [30] commonly only evaluated on short tracking challenges like WildTrack. From 1,000 frames to the full AICity’24 length, MCBLT only drops by $4.58$ HOTA points compared to the massive performance decrease of $43.35$ utilizing only a Kalman filter. Such trackers fail to tackle the long-term tracking challenges common in MTMC benchmarks.
Long-term Asc. Max. Window HOTA AssA DetA
Heuristics [5] 480 51.87 31.26 86.12
960 62.50 45.25 86.36
1920 71.88 59.78 86.47
3840 76.80 68.22 86.49
Model (Ours) Global 81.22 76.19 86.94
Table 5: SUSHI inference ablation on the AICity’24 test set. Rows 1 to 4 rely on heuristic matching to track overlapping sliding windows inferred by the SUSHI GNN hierarchy. Our final tracking solution (the last row) applies a global merging block to associate without graph overlaps.
Method # of frames HOTA AssA DetA
MCBLT 1,000 95.09 95.77 94.41
3,000 93.62 92.37 94.89
5,000 91.85 88.82 94.98
10,000 91.06 88.33 93.89
23,994 90.51 88.05 93.03
BEV-KF (baseline) 1,000 63.49 45.29 89.01
3,000 45.28 22.86 89.71
5,000 35.54 14.23 88.79
10,000 27.91 8.87 87.80
23,994 20.15 4.67 86.91
@@ 5 Conclusion
This paper proposed MCBLT, an accurate and robust multi-camera 3D detection and tracking framework, based on early multi-view aggregation, in environments monitored by static multi-camera systems. The proposed framework has impressive generalizability for diverse scenes and camera settings. It also has a powerful capability for long-term association to track objects in very long videos. As a result, MCBLT achieves SOTA on both the AICity’24 dataset and the WildTrack dataset.
@@ B.3 Improvements with Detection Association
We compared the tracking performance of MCBLT with and without the proposed 2D-3D detection association algorithm in Tab. 9. The baseline result is based on the ReID features extracted from the large projected 3D bounding boxes shown in Fig. 3. With the noisy background or other objects included in the image crops, ReID feature quality will be significantly affected.
@@ Appendix C ReID Feature Quality Analysis
We evaluated the ReID feature quality by the mean average precision (mAP), rank-1, rank-5, and rank-10 accuracies. The evaluation results are shown in Tab. 10. We found that the feature quality on the WildTrack dataset is worse than that on the AICity’24 dataset. This is because i) WildTrack is a real-world dataset with more noises and diverse illuminations from different camera views; ii) 2D bounding box annotations are not as accurate as the synthetic AICity’24 dataset. Nevertheless, our MCBLT achieved impressive results on WildTrack based on these ReID features.
Appendix D Model Time Complexity Analysis
Although MTMC detection and tracking tasks do not usually require real-time performance and are tolerant to time delays, we record the running time of the proposed MCBLT pipeline in Tab. 11 to provide a rough impression of the complexity of the model. The model inference is conducted on one single NVIDIA A100 GPU, with 10 cameras in the scene. Our method achieves around 1.5 FPS end-to-end before any further model optimization. The 2D detection, ReID, and tracking models are very efficient and can operate in parallel with BEVFormer so that their running time is negligible.
Detection Tracking
BEVFormer   1.6 FPS SUSHI 452.7 FPS
DINO  65.0 FPS SOLIDER  58.9 FPS
Table 11: MCBLT model efficiency analysis.
Appendix E Overall Visualization
We also visualized MTMC detection and tracking results of MCBLT on the AICity’24 and WildTrack datasets. Please find the demos in the attachment.